# example.bib

### v1

@@ I Introduction
In vision-based robotic manipulation, end-to-end (E2E) visuomotor policies which map pixel inputs directly to control signals are gaining popularity due to their ability to model complex robot behaviors and motion modalities [levine2016end, pmlrv87kalashnikov18a, pmlrv155zeng21a, pmlrv205shridhar23a]. However, training policies from scratch is restricted to simulation environments or requires large human demonstration datasets which are difficult to collect and scale [robomimic2021, mandlekar2023mimicgen].
In computer vision, this low data issue is tackled by pretraining a universal visual encoder on large image datasets and adopting it for diverse downstream tasks [Mahajan2018ECCV, rich2014cvpr]. Recent advancements in robotic manipulation have mirrored this trend [nair2022rm, Xiao2022, ma2022vip, parisi2022the, radosavovic2022realworld]. Specifically, they pretrain a visual encoder using out-of-domain (OOD) natural image and video datasets, such as ImageNet [imagenetcvpr09] and Ego4D [grauman2022ego4d], capturing semantic information [ma2022vip] or temporal dynamics [nair2022rm] in images or videos. The trained encoders are subsequently used frozen (without weights update) to extract visual features when training policies across various manipulation tasks. Significant performance improvements over E2E training are reported across diverse policy and task domains, particularly in low-data regimes (5 to 100 demonstrations).
@@
Figure 1: Formulation of Visuomotor Policies. Although in E2E models, a visuomotor policy can be viewed as a visual encoder $f$ followed by a regression head $\psi$. The two components as a whole are treated as the policy. In OOD pretraining, a functional separation of the visual encoder emerges and only the regression head acts as the policy.
@@
However, the validity of this functional separation and the potential benefits of pretraining is compromised if its performance margin cannot be maintained under certain circumstances. In this paper, we establish a performance benchmark comparing OOD-pretrained encoders with the state-of-the-art performance achieved by E2E policies [NIPS1988812b4ba2, robomimic2021, chi2023diffusionpolicy]. Our benchmarking results (see Tab. III) reveal that policies utilizing frozen pretrained encoders lose their performance advantage and result in an average success rate gap of 42% in simulation.
To further investigate the causes of the performance gap, given the difference in policy formulation between the two training paradigms, we hypothesize that the performance gap arises because E2E-trained visual encoders play an active role in decision-making, a capability OOD-pretrained encoders do not develop. To test this hypothesis, we introduce Visual Alignment Testing (VAT), an experimental framework designed to quantitatively assess the impact of visual encoders on decision-making (detailed in Section III.C). Additionally, we employ Fullgrad saliency maps [Srinivas2019FullGradientRF] to qualitatively demonstrate how visual encoders contribute to decision-making through their pixel-wise activations. Experiments were conducted in both simulations and the real world.
Although the hypothesis may seem straightforward within the E2E paradigm, it has not yet been explicitly identified and verified. Additionally, its validity is becoming increasingly important for the community as OOD pretraining is built upon the assumption of the functional separation. This work provides initial quantitative and qualitative evidence explaining the underperformance of OOD pretraining for robotic manipulation, underscoring the overlooked role of visual encoders in visuomotor policies. The main contributions are:
@@
  • Our results prompt a reflection on the role of visual encoders in visuomotor policies and the validity of their functional separation in pretraining, offering valuable insights for future research.
@@ II Related Work
Recent works have explored developing a universal encoder for manipulation which is trained on OOD datasets and leveraging the encoder frozen for policy regression [radosavovic2023robot, Yao2022MaskedCR, ChaneSane2023LearningVP, Wu2023PolicyPF]. PVR [parisi2022the] finds that encoders pretrained on computer vision tasks can be transferred to behavior cloning. MVP [Xiao2022], R3M [nair2022rm] and VIP [ma2022vip] investigate into self-supervised representation learning for control. Specifically, MVP trains a masked autoencoder (MAE) [He2021MaskedAA] on natural images as the visual encoder. Radosavovic et al. [radosavovic2022realworld] extend this idea from simulation to real robot imitation learning. R3M uses a combination of time-contrastive learning and video-language alignment to capture environmental dynamics and semantics. VIP generates representations and dense visual rewards by reforming it into a value function optimization problem.
Several endeavours were made to benchmark the advances in OOD-pretrained encoders for manipulation. Jing et al. [jing2022exploring] explore three components in pretraining: datasets, models and training objectives. A more detailed dataset-centric analysis for pre-training is proposed by [dasari2023an]. Hu et al. [Yingdong2023for] conduct empirical evaluations for diverse pretrained vision models across different policies. However, these benchmarking studies primarily focus on comparing the performance of different pretrained encoders, without providing a direct comparison with E2E baselines. Hansen et al. [hansen2022on] show a simple Learning-from-Scratch baseline is competitive with state-of-the-art pretrained models. However, they simply attribute the performance gap between pretrained and E2E models to domain gaps, without offering an in-depth understanding of the underlying reasons. Additionally, the role of visual encoders in visuomotor policies remains underexplored. This paper takes steps towards addressing these gaps by establishing a performance benchmark between pretrained encoders and their E2E-trained counterparts and facilitating a deeper understanding on the role of visual encoders with both quantitative and qualitative evidence.
@@ III Methodology
Method Objective Model N-o-P (M) Dataset Supervision
ImageNet Cross Entropy ResNet18 11 ImageNet Class
R3M Temporal Contrastive ResNet18 11 Ego4D SSL
MVP MAE ViT-S 22 Multi SSL
VIP Value Optimization ResNet50 23 Ego4D SSL
TABLE I: Overview of Pretrained Encoders. We summarize the pretrained encoders considered in the benchmark comparison in terms of training objectives, model architectures, numbers of parameters (in millions), training datasets and types of supervision. MVP uses a combination of Ego4D, ImageNet and HoI [Xiao2022] datasets for the training. SSL stands for self-supervised learning.
@@ III-A Preliminary
Learning-based robotic manipulation commonly develops visuomotor policies using reinforcement learning (RL) [kober2013reinforcement] or imitation learning (IL) [hussein2017imitation]. In this paper, we primarily focus on IL. An IL policy learns action distribution from expert demonstration data $(\mathcal{O},\mathcal{P},\mathcal{A})$, including image observation $\mathcal{O}$, proprioceptive state $\mathcal{P}$ and robot action $\mathcal{A}$. Under the E2E-training scenario, given an observation state $o_{t}\in\mathcal{O}$ and a proprioceptive state $p_{t}\in\mathcal{P}$, a policy $\pi$ directly maps the states to robot actions $\hat{a}_{t}=\pi(o_{t},p_{t})$. The policy network $\pi$ can be viewed as two parts: a visual encoder $f_{\theta}$ which extracts visual features from $o_{t}$ and a regression head $\psi_{\phi}$ which predicts the actions, $\hat{a}_{t}=\psi_{\phi}(f_{\theta}(o_{t}),p_{t})$. The weights $\theta\in\mathbb{R}^{n_{1}}$ and $\phi\in\mathbb{R}^{n_{2}}$ are updated simultaneously to minimize the objective $||a_{t}-\hat{a}_{t}||_{i},\;\text{for}\;i=1,2$. For OOD pretraining, an OOD-pretrained visual encoder $f^{*}$ is applied frozen to image inputs and only $\phi$ is updated during training. We organize the rest of this section by introducing benchmark setting, design of Visual Alignment Testing (VAT) and the Fullgrad saliency map.
Figure 2: Visual Alignment Testing. The experimental framework we proposed provides quantitative evidence that E2E-trained visual encoders play an active role in decision-making, detailed in Section III.C.
@@
TABLE II: Task Summary. ActD: robot action dimension, ObsD: image observation dimension, Data: number of demonstrations, Step: max rollout steps in simulation during evaluation, RichC: whether the task requires contact-rich manipulation, LongH: whether the task requires learning multiple behaviors together. Tasks for benchmarking are marked in blue, while the others are used in VAT.
III-B Benchmark Setting
III-B1 Pretrained Encoders
We compare the performance of three state-of-the-art pretrained encoders specifically designed for control: R3M [nair2022rm], MVP [radosavovic2022realworld] and VIP [ma2022vip]. An ImageNet pretrained ResNet18 [he2015deep] model is also considered due to its widespread use in computer vision. Tab. I shows an overview of the pretrained encoders. During the policy training phase, all encoders are frozen, adhering to the original evaluation protocols without modifications. For those with multiple variants, we adopt the smallest version of publicly available models to ensure a fair comparison. The pretrained encoders are evaluated in combination with various regression heads, as detailed in Section III.B.2.
@@ III-C Visual Alignment Testing
We hypothesise that E2E-trained visual encoders are playing an active role in decision-making. This hypothesis, if shown to hold, may explain the underperformance of pretrained encoders which lack task data supervision due to the functional separation in OOD pretraining. However, the implicit nature of neural networks presents a challenge in directly validating the hypothesis. Therefore, we propose Visual Alignment Testing, a proof-by-contradiction approach as outlined below.
@@
where $d$ is the metric for policy performance, and $\mathbb{E}$ stands for the expectation of testing rollouts.
@@
A contradiction occurs. This would imply that the assumption of visual encoders not playing an active role in decision-making is proven to be incorrect. A threshold $\delta$ is introduced to capture training randomness.
@@
where $N$ is the number of testing environments, $\hat{\mu}_{D}$ is the estimated mean value of $D_{i}$.
@@
We reject $H_{0}$ if $t$ is greater than the critical value given a preset significance level $\alpha\in(0,1)$.
@@ IV-A2 Implementation Detail
(a) Mug $\rightarrow$ Spam
(b) $\textit{Square}^{\star}\rightarrow$ Nut
(c) $\textit{Can}^{\star}\rightarrow$ Pick Place
(d) Reach $\rightarrow$ PnP
Figure 3: Tasks for Visual Alignment Testing. We conduct VAT on three groups of tasks in simulation and one in the real world, illustrated in the order of $\mathcal{T}_{0}\rightarrow\mathcal{T}_{1}$.
@@ IV-B Experimental Result
(c) Multi-task Encoder
@@
[Result 2] The performance gap arises because E2E-trained visual encoders play an active role in decision-making.
@@
Qualitative Result. To understand how visual encoders contribute to decision-making, we present the saliency maps of encoders $f_{1}$ and $f_{0}$ trained on VAT tasks in Fig. 5. Notably, E2E-trained encoders are primarily activated by specific task-relevant areas. For example, given the same input image of Spam, E2E-trained $f_{1}$ mainly focuses on the box of spam the robot is reaching, while $f_{0}$ (trained on Mug) focuses on the mugs. Despite the identical visual content in the two task inputs, the encoders, updated by different motor data, develop preferences that reflect the task intent. Comparing the saliency maps of E2E and OOD-pretrained encoders in Fig. 5, the underperforming of pretrained encoders might come from their failure to effectively focus on task-related areas. For instance, in Square, the E2E model activates on the center of the square nut and the rod, whereas Imagenet, R3M and VIP do not give sufficient attention to the task-relevant objects.
In Fig. 5, we also demonstrate that the decision-making ability of visual encoders may extend to multi-task settings, where encoders differentiate between tasks and focus on task-relevant features accordingly. Additionally, as shown in Fig. 5, ImageNet features recognize large objects in a scene, capturing most key information. This property of non-specific yet information-complete explains why ImageNet performs relatively better than other pretrained encoders.
@@ V Conclusion and Limitation
In this paper, we identify and analyze an underexplored question arising from the introduction of OOD pretraining to visuomotor policy learning: E2E-trained visual encoders actively contribute to the decision-making process by focusing on specific task-relevant areas in the visual inputs. Both quantitative evidence derived from the proposed Visual Alignment Testing, applied in simulations and the real world, and qualitative results provided by FullGrad saliency maps support this claim. This finding provides an initial explanation for the underperformance of OOD pretrained encoders compared to E2E training, as demonstrated in our benchmarking. Our results indicate that future works may benefit from training visual encoders on manipulation datasets that span diverse task domains and incorporating task conditions and context information into the encoders.
One of the limitations of this work is that we did not consider very low data regimes addressed in some pretraining works, e.g. working with 5 demonstrations. Decreasing the training data size dramatically is likely to negatively affect the decision-making ability of E2E visual encoders. An investigation of smaller dataset sizes and the exact threshold where this impact becomes significant is however an interesting direction for future work.

### v2

@@ I Introduction
In vision-based robotic manipulation, end-to-end (E2E) visuomotor policies which map pixel inputs directly to control signals are gaining popularity due to their ability to model complex robot behaviors and motion modalities [levine2016end, pmlrv87kalashnikov18a, pmlrv155zeng21a, pmlrv205shridhar23a]. However, training policies from scratch is still restricted to simulation environments or requires large human demonstration datasets which are difficult to collect and scale [robomimic2021, mandlekar2023mimicgen].
In computer vision (CV), the limited availability of training data is addressed by pre-training a universal visual encoder on large image datasets and adopting it for diverse downstream tasks [Mahajan2018ECCV, rich2014cvpr]. Recent advances in robotic manipulation have mirrored this trend [nair2022rm, Xiao2022, ma2022vip, parisi2022the, radosavovic2022realworld]. Specifically, they pretrain a visual encoder using out-of-domain (OOD) natural image and video datasets, such as ImageNet [imagenetcvpr09] and Ego4D [grauman2022ego4d], capturing semantic information [ma2022vip] or temporal dynamics [nair2022rm] in images or videos. Subsequently, the trained encoders are used frozen (weight updates are prohibited) to extract visual features when training policies across various manipulation tasks. Performance improvements over E2E training are reported across diverse policy and task domains, particularly in low-data regimes (5 to 100 demonstrations).
@@
Figure 1: Formulation of Visuomotor Policies. In E2E formulation, although a visuomotor policy can be viewed as a visual encoder $f$ followed by a regression head $\psi$, the two components as a whole are treated as the policy. In OOD pretraining, a functional separation of the visual encoder emerges and only the regression head acts as a policy.
@@
However, the validity of this functional separation and the potential benefits of pretraining is compromised if its performance margin cannot be maintained under certain circumstances. In this paper, we establish a performance benchmark comparing models with OOD-pretrained encoders with the state-of-the-art performance achieved by E2E policies [NIPS1988812b4ba2, robomimic2021, chi2023diffusionpolicy]. Our benchmarking results (see Tab. III) reveal that policies utilizing frozen pretrained encoders lose their performance advantage, with an average success rate gap of 42% in simulation.
To further investigate the causes of the observed gap and inspired by the difference in policy formulation between the two paradigms, we hypothesize that the performance gap arises because E2E-trained visual encoders play an active role in decision-making, a capability OOD-pretrained encoders do not develop. To verify this hypothesis, we introduce Visual Alignment Testing (VAT), an experimental framework designed to quantitatively assess the impact of visual encoders on decision-making (detailed in Section III.C). Additionally, we employ Fullgrad saliency maps [Srinivas2019FullGradientRF] to qualitatively demonstrate how visual encoders contribute to decision-making through their pixel-wise activations on the visual inputs. Experiments were conducted in both simulations and the real world.
Although the hypothesis may seem straightforward, it has to the best of our knowledge not previously been explicitly identified and verified. Additionally, its validity is becoming increasingly important for the community as OOD pretraining is built upon the assumption of the functional separation, and reports of pretraining underperformance are growing [liu2023libero, hansen2022on, sharmalossless, lin2023spawnnet]. This work provides initial quantitative and qualitative evidence that may in part explain the underperformance of OOD pretraining for robotic manipulation, underscoring the overlooked role of visual encoders in visuomotor policies. The main contributions are:
@@
  • Our results prompt a reflection on the role of visual encoders in visuomotor policies and the validity of their functional separation in pretraining that may be relevant for developing future methods in this area.
@@ II Related Work
Recent works have explored developing a universal encoder for manipulation which is trained on OOD datasets and leverages the frozen encoder for policy regression [radosavovic2023robot, Yao2022MaskedCR, ChaneSane2023LearningVP, Wu2023PolicyPF, zeng2024learning, karamcheti2023voltron]. PVR [parisi2022the] find that encoders pretrained on computer vision tasks can be transferred to behavior cloning. MVP [Xiao2022], R3M [nair2022rm] and VIP [ma2022vip] investigate self-supervised representation learning for control. Specifically, MVP trains a masked autoencoder (MAE) [He2021MaskedAA] on natural images as the visual encoder. Radosavovic et al. [radosavovic2022realworld] extend this idea from simulation to real robot imitation learning. R3M uses a combination of time-contrastive learning and video-language alignment to capture environmental dynamics and semantics. VIP generates representations and dense visual rewards by reforming it into a value function optimization problem.
Several efforts have been made to benchmark the advances in OOD-pretrained encoders for manipulation. Jing et al. [jing2022exploring] explore three components in pretraining: datasets, models and training objectives. A more detailed dataset-centric analysis for pre-training is proposed by [dasari2023an]. Hu et al. [Yingdong2023for] conduct empirical evaluations for various pre-trained vision models under different policies. However, these benchmarking studies primarily focus on comparing the performance of different pretrained encoders, without considering E2E-trained baselines. Hansen et al. [hansen2022on] show a simple Learning-from-Scratch baseline is competitive with state-of-the-art pretrained models. However, the authors attribute the performance gap between pretrained and E2E models to domain gaps, without a deeper investigation into the underlying causes. Additionally, the role of visual encoders in visuomotor policies remains underexplored. This paper takes steps towards addressing these gaps by benchmarking pretrained encoders against their E2E counterparts and providing quantitative and qualitative evidence into the role of visual encoders.
@@ III Methodology
Method Objective Model N-o-P (M) Dataset
ImageNet Cross Entropy ResNet18 11 ImageNet
R3M Temporal Contrastive ResNet18 11 Ego4D
MVP MAE ViT-S 22 Multi
VIP Value Optimization ResNet50 23 Ego4D
TABLE I: Overview of Pretrained Encoders. We summarize the pretrained encoders considered in the benchmark comparison in terms of training objectives, model architectures, numbers of parameters (in millions), and training datasets. MVP uses a combination of the Ego4D, ImageNet and HoI [Xiao2022] datasets for the training.
@@ III-A Preliminary
Learning-based robotic manipulation commonly develops visuomotor policies using reinforcement learning (RL) [kober2013reinforcement] or imitation learning (IL) [hussein2017imitation]. In this paper, we primarily focus on IL. An IL policy learns action distribution from expert demonstration data $(\mathcal{O},\mathcal{P},\mathcal{A})$, including image observation $\mathcal{O}$, proprioceptive state $\mathcal{P}$ and robot action $\mathcal{A}$. Under the E2E-training scenario, given an observation state $o_{t}\in\mathcal{O}$ and a proprioceptive state $p_{t}\in\mathcal{P}$, a policy $\pi$ directly maps the states to robot actions $\hat{a}_{t}=\pi(o_{t},p_{t})$. The policy network $\pi$ can be viewed as two parts: a visual encoder $f_{\theta}$ which extracts visual features from $o_{t}$ and a regression head $\psi_{\phi}$ which predicts the actions, $\hat{a}_{t}=\psi_{\phi}(f_{\theta}(o_{t}),p_{t})$. The weights $\theta\in\mathbb{R}^{n_{1}}$ and $\phi\in\mathbb{R}^{n_{2}}$ are updated simultaneously to minimize the objective $||a_{t}-\hat{a}_{t}||_{i},\;\text{for}\;i=1,2$. For OOD pretraining, a pretrained visual encoder $f^{*}$ is applied frozen to image inputs and only $\phi$ is updated during training.
III-B Benchmark Setting
III-B1 Pretrained Encoders
We evaluate the performance of three state-of-the-art pretrained encoders designed for control: R3M [nair2022rm], MVP [radosavovic2022realworld], and VIP [ma2022vip]. Additionally, we include an ImageNet-pretrained ResNet18 [he2015deep] due to its widespread use in CV. Other popular pretrained CV models are excluded, as their performance has already been extensively benchmarked in prior studies [Yingdong2023for]. During the policy training phase, all encoders are frozen, adhering to the original evaluation protocols without modifications. For models with multiple variants, we adopt the smallest publicly available version to ensure a fair comparison. The pretrained encoders are evaluated in combination with various regression heads, as detailed in Section III.B.2.
Figure 2: Visual Alignment Testing. The experimental framework we propose provides quantitative evidence that E2E-trained visual encoders play an active role in decision-making, detailed in Section III.C.
@@ III-B1 Pretrained Encoders
TABLE II: Task Summary. ActD: robot action dimension, ObsD: image observation dimension, Data: number of demonstrations, Step: max rollout steps in simulation during evaluation, RichC: whether the task requires contact-rich manipulation, LongH: whether the task requires learning multiple behaviors together. Tasks for benchmarking are marked in blue, while the others are used in VAT, see Fig. 3.
@@ III-B3 Testing Environments
(a) Mug $\rightarrow$ Spam
(b) $\textit{Square}^{\star}\rightarrow$ Nut
(c) $\textit{Can}^{\star}\rightarrow$ Pick Place
(d) Reach $\rightarrow$ PnP
Figure 3: Tasks for Visual Alignment Testing. We conduct VAT on three groups of tasks in simulation and one in the real world, illustrated in the order of $\mathcal{T}_{0}\rightarrow\mathcal{T}_{1}$, see Sec. III.C.
@@ III-C Visual Alignment Testing
We hypothesise that E2E-trained visual encoders are playing an active role in decision-making. This hypothesis, if shown to hold, may explain the underperformance of OOD pretraining which lacks task data supervision due to the inherent functional separation. However, the implicit nature of neural networks makes direct validation challenging. Therefore, we propose Visual Alignment Testing, a proof-by-contradiction approach outlined below.
@@
where $d$ is the metric for policy performance, and $\mathbb{E}$ stands for the expectation under testing rollouts.
@@
A contradiction occurs. This would imply that the assumption of visual encoders not playing an active role in decision-making is shown to be incorrect. A threshold $\delta$ is introduced to capture training randomness.
@@
where $N$ is the number of testing rollouts, $\hat{\mu}_{D}$ is the estimated mean value of $D_{i}$.
@@
We reject $H_{0}$ if $t$ is greater than the critical value given a preset significance level $\alpha\in(0,1)$:
@@ IV-B Experimental Result
(c) A Multi-task Encoder
@@
[Result 2] The performance gap stems from E2E-trained visual encoders actively influencing decision-making.
@@
Qualitative Result. To understand how visual encoders contribute to decision-making, we present the saliency maps of encoders $f_{1}$ and $f_{0}$ trained on VAT tasks in Fig. 5. Notably, E2E-trained encoders are primarily activated by specific task-relevant areas. For example, given the same input image of Spam, E2E-trained $f_{1}$ mainly focuses on the box of spam the robot is reaching, while $f_{0}$ (trained on Mug) focuses on the mugs. Despite identical visual inputs, encoders updated with different motor data develop preferences that align with task intent. Comparing the saliency maps of E2E and OOD-pretrained encoders in Fig. 5, the underperforming of pretrained encoders might come from their failure to focus on task-related areas effectively. For instance, in Square, the E2E model activates on the center of the square nut and the rod, whereas Imagenet, R3M and VIP do not give sufficient attention to the task-relevant objects.
Encoder E2E Pretrain Pretrain+Finetune
Succ. rate 0.68 0.26 0.61
TABLE V: Finetuning Results. All finetuned models are trained and evaluated following the same protocol in Section IV.A. All models use BC as the regression head and averaged over PushT, Can and Square. The results of pretrained encoders are averaged over ImageNet, R3M and VIP.
In Fig. 5, we demonstrate that the decision-making ability of E2E encoders extends to multi-task settings, where they differentiate between tasks and focus on task-relevant features accordingly. This alleviates concerns that their decision-making ability results from overfitting to a single task. Furthermore, Fig. 5 shows that ImageNet features effectively recognize large objects in a scene, capturing most key information. This non-specific yet information-rich property likely explains ImageNet’s superior performance compared to other pretrained encoders. For completeness, Tab. V shows that finetuning pretrained encoders can enhance their performance, which does not contradict our conclusions. Investigating optimal finetuning methods and how they influence the representations lies beyond the scope of this study.
@@ V Conclusion and Limitation
In this paper, we identify and analyze an underexplored question arising from the introduction of OOD pretraining to visuomotor policy learning: E2E-trained visual encoders actively contribute to the decision-making process by focusing on specific task-relevant areas in the visual inputs. Both quantitative evidence derived from the proposed Visual Alignment Testing, applied in simulations and the real world, and qualitative results provided by FullGrad saliency maps support this claim. This finding provides an initial explanation for the underperformance of OOD pretrained encoders, as demonstrated in the benchmarking results. Our results indicate that future works may benefit from training visual encoders on manipulation datasets that span diverse task domains and incorporating task conditions and context information into the encoders.
One limitation of this work is the exclusion of very low-data regimes explored in some pretraining studies, e.g. working with 5 demonstrations. A significant reduction in training data is likely to impair the decision-making ability of the E2E visual encoders. Investigating the effects of smaller dataset sizes and identifying the threshold at which performance degradation becomes significant is an interesting avenue for future research. Additionally, our analysis is confined to the encoder level. Examining how the policy further leverages these representations is another direction worth exploring.